One more than 4 times a number is 13. What is the number?
Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

let x be the number
4x+1=13
4x=13-1
4x=12
/4
x=3

Answer by rfer(16322)

4x+1=13
4x=12
x=3